# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v2

@@ 7.3 Practical Examples
To demonstrate the practical application of ICE-PIXIU, we have provided examples of model response as shown in Appendix Table 12. It can be concluded that ICE-INTERN provides more accurate responses in tasks such as FinED, FinNC, and FinRE, and offers more comprehensive responses in tasks like FinER and FinQA. Particularly in FinER, ICE-INTERN is able to provide all entity names and their corresponding types. Overall, ICE-INTERN outperforms in accuracy and comprehensiveness in financial question answering.
8 Limitations
@@
While this study has positive aspects, we acknowledge limitations: 1) Parameter Constraints: Limited resources hinder training models with over 7B parameters, impacting performance on large datasets. 2) Backbone LLM Adaptability: Fine-tuned model performance varies based on backbone model adaptability to specific tasks, requiring task-specific selection. 3) Prompt Quality Variability: Even with expert annotation, instruction prompts may perform differently across models, affecting model assessment. 4) Risk of Misuse: Open-source models facilitate research, but commercial misuse can lead to financial risks, necessitating effective regulation.
9 Conclusion
In conclusion, we presented innovative cross-lingual ICE-PIXIU financial framework, which consists of the pioneering cross-lingual bilingual financial LLM, ICE-INTERN, and the the evaluation benchmark, ICE-FLARE. The bilingual capability of ICE-INTERN empowers global finance by breaking language barriers, while ICE-FLARE enables comprehensive cross-lingual assessments for various financial specific tasks. The open access nature of ICE-PIXIU promotes collaboration and research in financial NLP. The framework underscores the significance of different types of instruction data in enhancing performance. Moreover, the framework provides a unified solution for diverse financial applications and ensures reliable cross-lingual consistency through rigorous evaluation. Overall, ICE-PIXIU contributes to the progress of research and development in financial NLP, fostering innovation within the field.
Acknowledgement
We would also like to acknowledge the financial support received from the General Program of Applied Basic Research Programs of Science and Technology Department of Yunnan Province (No.202301AT070184), the Science Research Fund Project of Education Department of Yunnan Province (No.2023J0019), the Key R&D Project of Hubei Province (No.2021BAA029) and the Open Project Program of Yunnan Key Laboratory of Intelligent Systems and Computing (No.ISC22Y08).
@@ Acknowledgement
A Examples of Context and Answer in Financial Task
Table 4-6 presents the examples of context description and annotated answer in various financial task in Chinese, English and translation datatsets.
@@ A Examples of Context and Answer in Financial Task
B Details Information of Raw Data Statistics
Table 7 presents the specifics of the raw data used to construct the Chinese-English bilingual dataset for multi-task financial instruction tuning.
| Lang | D_T | NLP_T | S_T | Dataset | Raw | Instruction | Evaluation | Data Source | License |
| StockB | 9,812 | 9,812 | 1,962 | social texts | Apache-2.0 | ||||
| FinSM | BQC | 120,000 | 110,000 | 10,000 | bank service logs | Public | |||
| AFQMC | 38,650 | 38,650 | 4,316 | online chat service | Apache-2.0 | ||||
| FinNC | NL | 7,955 | 7,955 | 884 | news articles | Public | |||
| NL2 | 7,955 | 7,955 | 884 | news articles | Public | ||||
| FinNJ | NSP | 4,499 | 4,499 | 500 | social texts | Public | |||
| FinAS | FinevalF | 1,115 | 1,115 | 222 | financial exam | Apache-2.0 | |||
| FinRE | RE | 14,973 | 14,973 | 1,489 | news, entity pairs | Public | |||
| FinER | CNER | 1,685 | 1,685 | 337 | financial reports | Public | |||
| 20CCKS | 372,810 | 45,796 | 9,159 | news, reports | CC BY-SA 4.0 | ||||
| 21CCKS | 8,000 | 7,000 | 1,400 | news, reports | CC BY-SA 4.0 | ||||
| 22CCKS | 109,555 | 59,143 | 11,829 | news, reports | CC BY-SA 4.0 | ||||
| FinQA | CEnQA | 8,281 | 668 | 133 | earnings reports | MIT license | |||
| CConFinQA | 12,594 | 1,189 | 237 | earnings reports | MIT license | ||||
| FiQA-SA | 1,173 | 1,173 | 235 | news headlines,tweets | Public | ||||
| FinCC | German | 1,000 | 1,000 | 200 | credit records | CC BY-SA 4.0 | |||
| Australian | 690 | 690 | 139 | credit records | CC BY-SA 4.0 | ||||
| ConFinQA | 3,458 | 12,594 | 1,490 | earnings reports | MIT License | ||||
| DOF | EN-DOF | FinER | FINER-ORD | 1,075 | - | 1,075 | news articles | CC BY-SA 4.0 | |
| FinTS | ECTSUM | 495 | - | 495 | earning call transcipts | Public | |||
| EDTSUM | 2,000 | - | 2,000 | news articles | Public | ||||
| FinDC | FOMC | 496 | - | 496 | FOMC transcripts | CC BY-SA 4.0 |
C Example of Chinese Dataset Translation
Figure 5 presents the example of translating BigData22 into Chinese CBigData22 Chinese dataset using human-written specific translation prompts.
D Human Evaluation for Expert-written Prompts
Table 8-9 presents the example of human evaluation for domain-expert written prompts on various financial specific task
| Specific | Prompt | Input | Target | Response | Metric | |||
| Task | acc | nat | inf | AVG | ||||
| Sentiment Analysis (FinSA) | 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。 | 涨空间异常巨大,中铁装配,实在是未来的大牛股啊 | 积极 | 积极 | 3 | 3 | 3 | 3 |
@@ D Human Evaluation for Expert-written Prompts
| Specific | Prompt | Input | Target | Response | Metric | |||
| Task | acc | nat | inf | AVG | ||||
@@
E Examples of Instruction Prompt in Chinese Task
Table 10 presents the example of the Chinese prompt and corresponding English translation for each specific financial task.
| Specific Task | Chinese Prompt | English Translation |
| Sentiment Analysis (FinSA) | 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。 | Determine the sentiment of the sentences in the provided financial news article, identifying whether the sentiment is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’. |
| Sentiment Matching (FinSM) | 判断给出的两个金融文本表达的意思是否相似,你只需要回答是或否。现在判断下面两个句子是否相似: | Determine whether the two provided financial texts convey similar meanings; you only need to answer ’yes’ or ’no’. Now determine if the following two sentences are similar. |
| Negative Judgment (FinNJ) | 根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’ | Based on the given financial news and entities, you need to determine whether the given entities contain negative messages. You only need to simply answer ’yes’ or ’no’. |
| Question Answering (FinQA) | 你需要分析金融文本,根据内容回答相关问题。如果你对答案感到迷茫,可以回答’无相应参数’。 | You need to analyze financial texts and answer related questions based on the content. If you feel confused about the answer, you can reply with ’No relevant parameters’. |
F Evaluation Results of LLMs, FinLLMs and ICE-INTERN variants
Table 11 presents the detailed evaluation results of different 6B-7B LLMs, baseline FinLLMs and variants of ICE-INTERN on the ICE-FLARE benchmark.
@@ F Evaluation Results of LLMs, FinLLMs and ICE-INTERN variants
ICE-INTERN-dlc-7B |
ICE-INTERN-dle-7B |
ICE-INTERN-dtt-7B |
ICE-INTERN-full-7B |
@@
| DOF | EN-OFT | FinER | Entity F1 | FINER-ORD | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.021 | 0.000 | 0.280* | 0.770* | 0.000 | 0.000 | 0.008 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | |
@@
G Examples of Query Response among Various LLMs
Table 12 presents the examples of query input and response among various LLMs on some specific financial task.
| Specific Task | Query Input (Including Prompt Description) | Model Response | ||

### v3

@@ 4.2. Practical Examples
To demonstrate the practical utility of ICE-PIXIU, we provide examples of model responses of ICE-PIXIU and other baseline including GPT-4 and Baichuan-7B, as outlined in in Appendix Table 12. It can be concluded that ICE-INTERN provides more accurate responses in FinED, FinNC, and FinRE tasks, and offers more comprehensive responses in FinER and FinQA tasks. Particularly in FinER, ICE-INTERN is capable of providing all entity names and their corresponding types. Overall, ICE-INTERN demonstrates superior in accuracy and comprehensiveness in financial question answering.
5. Limitations
@@
While this study has positive aspects, we acknowledge limitations: 1) Parameter Constraints: The limited resources hinder training models with over 7B parameters, impacting performance on large datasets. 2) Backbone LLM Adaptability: The fine-tuned model performance varies based on the backbone model adaptability to specific tasks, requiring task-specific selection. 3) Prompt Quality Variability: Even with expert annotation, the quality of instruction prompts may perform vary across models, affecting model assessment. 4) Risk of Misuse: Open-source models facilitate research, but commercial misuse raises concerns about financial risks. Effective regulation is therefore necessary to mitigate these risks.
6. Conclusion
We introduce a Chinese-English bilingual financial framework ICE-PIXIU featuring ICE-INTERN LLM, ICE-FIND instruction data, and ICE-FLARE evaluation benchmark to address global financial cross-lingual disparities. ICE-INTERN’s bilingual capability empowers global finance by breaking language barriers, while ICE-FLARE enables comprehensive cross-lingual assessments for various financial tasks. Notably, ICE-INTERN demonstrates superior performance over existing SOTA LLMs, including GPT-4, in Chinese financial tasks by effectively leveraging strategic instruction tuning and a diverse dataset for cross-linguistic transfer. ICE-PIXIU underscores the significance of diverse instruction data types and cross-lingual data translation transfer in enhancing performance. By rigorously evaluating DOT datasets, ICE-PIXIU supports robustness exploration of FinLLMs and ensures consistent performance assessments. ICE-PIXIU’s open-access framework drives exploration in financial NLP for complex tasks and multilingual applications.
@@ 6. Conclusion
A. A Example of Translating BigData22 to English
Figure 3 presents the examples of our strategy of translating the BigData22 dataset into Chinese using human-written specific prompts
B. Examples of Instruction Prompts in Chinese Task
Table 5 displays examples of Chinese human-written prompts and their corresponding English translations for each specific task.
| Specific Task | Chinese Prompt | English Translation |
| Sentiment Analysis (FinSA) | 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的,你的回答应该是其中一个。 | Determine the mood of the sentences in the financial news articles provided and identify whether the mood is ’Positive’, ’Negative’ or ’Neutral’. Your answer should be one of them. |
| Sentiment Matching (FinSM) | 判断给出的两个金融文本表达的意思是否相似,你只需要回答是或否。 | To judge whether the two financial texts are similar in meaning, you only need to answer ’Yes’ or ’No’. |
| Negative Judgment (FinNJ) | 根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’ | Based on the given financial news and entities, you need to determine whether the given entities contain negative messages. You only need to answer ’Yes’ or ’No’. |
| Question Answering (FinQA) | 你需要分析金融文本,根据内容回答相关问题。如果没有对应的答案,可以回答’无相应参数’。 | You need to analyze the financial text and answer relevant questions according to the content. If there is no corresponding answer, you can answer ’No corresponding parameter’. |
C. Examples of Introduction for Each Financial Task
Table 6 to Table 8 showcase examples of contextual text descriptions and target answers in various financial application scenarios.
@@ C. Examples of Introduction for Each Financial Task
D. Human Evaluation for Expert-written Prompts
Table 9and Table 10 illustrate examples of human evaluations for expert-written prompts across various financial-specific tasks.
| Specific Task | Prompt | Input | Target | Response | Metric | |||
| AC | NA | IN | AVG | |||||
| Sentiment Analysis (FinSA) | 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。 | 上涨空间异常巨大,中铁装配实在是未来的大牛股啊 | 积极 | 积极 | 3 | 3 | 3 | 3 |
@@ D. Human Evaluation for Expert-written Prompts
| Specific Task | Prompt | Input | Target | Response | Metric | |||
| AC | NA | IN | AVG | |||||
@@
E. The Detailed Evaluation Results of ICE-FLARE
Table 11 presents the detailed evaluation results of 6B-7B general LLMs, baseline FinLLMs, and our proposed ICE-INTERN variants.
@@ E. The Detailed Evaluation Results of ICE-FLARE
ICEdlc-7B |
ICEdle-7B |
ICEdtt-7B |
ICEfull-7B |
@@
| DOT | EN-DOT | FinER | Entity F1 | FINER-ORD | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.021 | 0.000 | 0.280* | 0.770* | 0.000 | 0.000 | 0.008 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | |
@@
F. Examples of Query Response in Various LLMs
Table 12 presents the examples of query and response among various LLMs on FinQA, FinTS, FinRE, FinED, FinER and FinNC task.
| Specific Task | Query Input | Model Response | ||